Need help solving theses with step by step will be on a test so I really want to know how to them do. Thanks.

Part A Determine the mass of CO2 produced by burning enough of methane to produce 2.25 x10 kJ of heat. CH,(g)-, 2 O2(g) → CO2(g) + 2 H2O(g) dMrsn =-802.3 kJ Express your answer using three significant figures. Submit G Give Up Part B Determine the mass of CO2 produced by burning enough of ℃,Hs(g1+ 502(g) → 3CO2(g) + 4H20(9) to produce 2.25 10 k.J of heat Express your answer using three significant figures. SubmitG Give Up Part C Determine the mass of CO octane to produce 2.25 x10 kJ of heat C,His (l), + 25/2 02(g) → 8 CO?(g) + 9 H20(g) by burning enough of Express your answer using three significant figures. SubmitG Give Up Part D Which fuel contributes least to global warming per k.J of heat produced? methane propane SubmitG Give Up
Answer #1

ANSWER FOR>>A)To produce 2.25*10^2 that is 225 kj of heat required 225/802.3 =0.28 moles of CH4.

When 1 mole CH4burn 1 mole of CO2 will produced.

Therefore 0.28 mole of CH4 burn ,0.28 mole of CO2 will produced.

1 mole of CO2 =44g of CO2.

Therefore 0.28 moles of CO2 =(44*0.28)g=12.3g of CO2.

MCO2=12.3g of CO2

ANSWER FOR >>B)To produced 225 kj of heat ,required 225/2217=0.10 moles of C3H8.

When 1 mole of C3H8 burn then 3 moles of CO2 will produced.

Therefore 0.10 moles of C3H8 burns then 0.10*3=0.30 moles of CO will produced.

1 mole CO2 =44 g of CO2.

Therefore 0.30 moles of CO2 =(0.30*44)g=13.02g og CO2.

MCO2=13.02g of CO2

ANSWER FOR >>C)To produced 225 kj of heat ,required=225/5074.1=0.044 moles of C8H18.

1 mole C8H18 burns then 8 moles of CO2 will produced.

Therefore 0.044 moles of C8H18 burns then 8*0.044=0.352 moles of CO2 will produced.

1 mole CO2=44g

Therefore 0.352 moles of CO2=(44*0.352)G=15.488g of CO2.

MCO2=15.488 g of CO2.

ANSWER FOR >>D)Since here we have find that more amount CH4 is required to produce 225 kj of heat ,that is here methane contributes least amount of global warming.